OpenStudy (anonymous):

Write the the slope intercept form of the equation of the line parallel to y=5x+2, which passes through the point (4,9)

OpenStudy (ivettef365):

what is the slope on equation y = 5x + 2

OpenStudy (anonymous):

5

OpenStudy (ivettef365):

good, so now use points (4,9), on equation y = 5x + b to find value of b

OpenStudy (anonymous):

9=5*4+b

OpenStudy (anonymous):

20-9=b right so far?

OpenStudy (ivettef365):

should be 9 - 20 = b

OpenStudy (anonymous):

ok b=-11

OpenStudy (ivettef365):

good, so you have your equation y = 5x -11

OpenStudy (anonymous):

y=5x-11?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

ok